OpenStudy (anonymous):

Solve the equation: x2 - 12x = -36

OpenStudy (anonymous):

First get it in this form then factor as usual x^2 - 12x + 36 = (x-6)(x-6)

OpenStudy (anonymous):

Do you see what I did?

OpenStudy (anonymous):

1) your first goal is to make the equation eqaul zero 2) add 36 to both sides,so the equation is x^2-12x+36 3) you automatically know that you have two negative roots since your middl term is negative when you add it and your lasst last term is positive when you multiply it 4) your final answer will be (x-6)(x-6) as a plus point, if you want to know what x equals to then do x-6=0, and x will equal 6 hope that helps :D